# A parallel framework for graphical optimal transport

### v1

@@ 2.2.1 Generalized Euler flow
One of the first applications of multi-marginal optimal transport, the generalized Euler flow (Benamou et al. 2015, Section 4.3), falls in fact into the category of a graph-structured problem. The Euler flow studies a relaxation of the incompressible Euler equation, where one seeks trajectories of particles between constrained initial and final data, that minimize the kinetic energy (Brenier 1989). Discretizing the problem into $J$ time points, we describe the flow by a $J$-mode multi-marginal optimal transport plan ${\mathbf{B}}\in{\mathbb{R}}_{+}^{d^{J}}$, where each marginal describes the density of the flow at time instance $j$, discretized into $d$ points. Note that the density of an incompressible flow, discretized uniformly over a given space, is described by a uniform distribution. Thus, we constrain each marginal of the transport plan by $P_{j}({\mathbf{B}})=\mu_{j}=\frac{1}{d}\boldsymbol{1}$. In the Euler flow, it is enforced that each particle at time 00 is transported to a given location at time $T$, given by a permutation $\sigma:\{1,\dots,d\}\to\{1,\dots,d\}$. This is modeled through a bi-marginal constraint $P_{1,J}({\mathbf{B}})=\boldsymbol{\mu}_{\{1,J\}}=\Pi_{\sigma}$, where $\Pi_{\sigma}\in{\mathbb{R}}^{d\times d}$ is the permutation matrix for the permutation $\sigma$. The Euler flow problem is thus formulated as the MOT problem
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}$ $\langle{\bf C},{\mathbf{B}}\rangle$ (12)
subject to $P_{j}({\mathbf{B}})=\mu_{j}\text{ ~{}~{}for }j=1,\dots,J;\quad P_{1,J}({\mathbf{B}})=\boldsymbol{\mu}_{\{1,J\}},$
@@
Benamou et al. (2015) also propose an alternative relaxed formulation, where the bi-marginal constraint in 12 is replaced by an additional cost term in 13, and the problem becomes
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j}\text{ for }j=1,\dots,J,$ (14)
@@
Both problems 12-13 and 14-15 can be described by a cyclic graph as illustrated in Fig. 1. In particular, we note that although the authors in Benamou et al. (2015) present only the computational method for the relaxed formulation 14-15, our proposed framework encompasses both formulations.
Figure 1: Illustration of graphical structures in some applications. Gray nodes correspond to fixed marginals, and white nodes are estimated in the problem. The dotted edge between $x_{1}$ and $x_{J}$ in (a) may represent either a bi-marginal constraint, as in 12-13, or a cost interaction, as in 14-15.
(a) Euler flow.
@@
(b) Wasserstein least squares.
@@ 2.2.2 Wasserstein least squares
$\min_{\pi,\mu_{0},\mu_{1}}$ $\int c(x_{0},x_{1},\cdots,x_{J},x_{J+1}){\rm{d}}\pi(x_{0},x_{1},\cdots,x_{J},x_{J+1})$ (17)
$P_{j}(\pi)=\rho_{t_{j}},~{}~{}j=1,2,\cdots,J.\qquad P_{0}(\pi)=\mu_{0},\qquad P_{J+1}(\pi)=\mu_{1}.$
By discretizing the integral in 17 to a summation, we can transform the problem into the form presented in 11. To enforce that the optimality of the coupling between $\mu_{0}$ and $\mu_{1}$, we can add a large penalty on it, leading to the modified cost
$c(x_{0},x_{1},\cdots,x_{J},x_{J+1})=\sum_{j=1}^{J}\|x_{j}-(1-t_{j})x_{0}-t_{i}x_{J+1}\|^{2}+\alpha\|x_{0}-x_{J+1}\|^{2}$ (18)
@@ 2.2.3 Wasserstein Splines
[…]
$\min_{\pi}\int c(x_{0},v_{0},\cdots,x_{J},v_{J}){\rm{d}}\pi(x_{0},v_{0},\cdots,x_{J},v_{J})\qquad\text{subject to }~{}~{}P_{x_{j}}(\pi)=\rho_{t_{j}},\quad j=0,1,\ldots,J$ (19)
@@
$\sum_{i=0}^{J-1}\frac{1}{t_{j+1}-t_{j}}\cdot\left[12\left\|\frac{x_{j+1}-x_{j}}{t_{j+1}-t_{j}}-v_{j}\right\|^{2}-12\left\langle\frac{x_{j+1}-x_{j}}{t_{j+1}-t_{j}}-v_{j},v_{j+1}-v_{j}\right\rangle+4\|v_{j+1}-v_{j}\|^{2}\right].$ (20)
@@ 3 Multi-marginal optimal transport with global regulatization and probabilistic graphical models
A line of recent works have drawn a connection between entropy-regularized multi-marginal optimal transport problems 4 with a graph-structured cost 10 and probabilistic graphical models (PGMs) (Haasler et al. 2021c, Fan et al. 2022a, Haasler et al. 2021b, Singh et al. 2022, Ringh et al. 2023, Zhou and Parno 2024). In this section we review and generalize these connections and the computational methods for solving these types of optimal transport problems.
3.1 A PGM formulation for multi-marginal optimal transport
A PGM provides a compact representation for the joint distribution of a collection of random vectors, which have dependencies that can be described by a graph (Koller and Friedman 2009). Define the tensors $K_{\alpha}({\bf x}_{\alpha})=\exp(-C_{\alpha}({\bf x}_{\alpha})/\epsilon)$ for $\alpha\in F$, where $C_{\alpha}$ are the cost tensors in 10, and $\epsilon$ is the regularization parameter in 4. Then, ${\bf K}$ in 7 has the form ${\bf K}({\bf x})=\prod_{\alpha\in F}K_{\alpha}({\bf x}_{\alpha})$. Moreover, to define the entropy regularization 5 for the graph-structured problem 11 we define the tensor ${\bf M}$ as ${\bf M}({\bf x})=\prod_{\alpha\in F_{\Gamma}}\boldsymbol{\mu}_{\alpha}({\bf x}_{\alpha})$. The factorizations of the tensors ${\bf K}$ and ${\bf M}$ can be used to define the joint probability distribution of a PGM,
$p({\bf x})=\frac{1}{Z}\left(\prod_{\alpha\in F}K_{\alpha}({\bf x}_{\alpha})\right)\left(\prod_{\alpha\in F_{\Gamma}}\boldsymbol{\mu}_{\alpha}({\bf x}_{\alpha})\right),$ (21)
where $Z$ is a normalization constant. Moreover, note that the solution tensor 6 can be written as ${\bf K}\odot{\bf U}\odot{\bf M}$ with ${\bf U}({\bf x})=\prod_{\alpha\in F_{\Gamma}}{\mathbf{u}}_{\alpha}({\bf x}_{\alpha})$. Thus, the scaling factors ${\mathbf{u}}_{\alpha}$, for $\alpha\in F_{\Gamma}$ can be interpreted as local potentials of the modified graphical model ${\bf K}({\bf x}){\bf U}({\bf x}){\bf M}({\bf x})$ The Iterative Scaling algorithm 1 finds potentials ${\mathbf{u}}_{\alpha}$, for $\alpha\in F_{\Gamma}$, such that the graphical model defined by the tensor ${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M}$ satisfies all the constraints $P_{\alpha}({\mathbf{B}})=\boldsymbol{\mu}_{\alpha}$, for $\alpha\in F_{\Gamma}$.
3.2 Iterative Scaling Belief Propagation
Each step in the Iterative Scaling algorithm involves calculating a projection $P_{j}({\bf K}\odot{\bf U}\odot{\bf M})$, which in terms of the PGM means inferring the distribution of the $j$-th variable. This Bayesian inference problem can be addressed by the Belief Propagation algorithm (Mezard and Montanari 2009, Part D), (Koller and Friedman 2009, Sec. 20.5.1). Belief propagation is known to converge when the underlying graph is a tree, i.e., does not have any loops, and reduces the complexity of computing the projections in the Iterative Scaling algorithm immensely. More precisely, whereas a brute force approach to compute the projections 3 scales exponentially, Belief Propagation scales only linearly in the number of marginals (Haasler et al. 2021c).
In Fan et al. (2022a), a similar approach has be utilized for general graphical models with possible loops, by partitioning the graph as a so-called junction tree. A junction tree (also called tree decomposition) describes a partitioning of a graph, where several nodes are clustered together, such that the interactions between the clusters can be described by a tree structure.
Definition 1 (Junction tree (Dawid 1992)).
@@ 3.2 Iterative Scaling Belief Propagation
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}{|{\bf M}})\qquad\text{ subject to }P_{c}({\mathbf{B}})=\boldsymbol{\mu}_{c},\text{ for }c\in\mathcal{C}_{\Gamma},$ (22)
@@
${\bf C}({\bf x})=\sum_{c\in\mathcal{C}}{\bf C}_{c}({\bf x}_{c}).$ (23)
Example 1.
Some examples of valid junction trees for our method are shown in Fig. 2 (for the Euler flow example in Fig. 1), Fig. 3 (for the Wasserstein least squares problem 17 in Fig. 1) , and Fig. 4 (for the Wasserstein splines problem in Fig. 1).
[…]
Figure 3: Junction tree for the graph for the Wasserstein least square problem from Fig. 1. The junction tree is constructed such that the leaves of the tree correspond to the constraints in the least square problem 17.
[…]
Figure 4: Junction tree of Wasserstein splines in Fig. 1.
while not converged do
$m_{i\to j}(s_{ij})$ $\leftarrow\sum_{x_{c_{i}}\setminus s_{ij}}{\bf K}_{c_{i}}({\bf x}_{c_{i}})\prod_{\ell\in N(i)\setminus j}m_{\ell\to i}(s_{\ell i}),\qquad c_{i}\in{\mathcal{C}}\setminus{\mathcal{C}}_{\Gamma}$ (24)
$m_{i\to j}(s_{ij})$ $\leftarrow\mu_{c_{i}}({\bf x}_{c_{i}})/m_{j\to i}({\bf x}_{c_{i}}),\qquad c_{i}\in{\mathcal{C}}_{\Gamma}.$ (25)
return $m_{i\to j}(s_{ij}),\ m_{j\to i}(s_{ji}),\ \forall(i,j)\in\mathcal{E}$
Similarly to the factorization 21, we can then write the probability distribution over the PGM following the junction tree decomposition $\mathcal{T}=(\mathcal{C},\mathcal{E})$ as a product
$p({\bf x})=\frac{1}{Z}\left(\prod_{c\in\mathcal{C}}{\bf K}_{c}({\bf x}_{c})\right)\left(\prod_{c\in\mathcal{C}_{\Gamma}}\boldsymbol{\mu}_{c}({\bf x}_{c})\right).$ (26)
Let $N(i)$ be the neighboring nodes of the node $i$. By extending the Iterative Scaling Belief Propagation methods in Haasler et al. (2021c), Fan et al. (2022a), we propose Algorithm 2 to solve 22. Upon convergence of Algorithm 2, the optimal transport plans on the clusters are given by
${\mathbf{B}}_{c}({\bf x}_{c})={\bf K}_{c}({\bf x}_{c})\prod_{\ell\in N(i)}m_{\ell\to i}(s_{\ell i}).$ (27)
The algorithm sends messages between neighboring cliques. The messages are thus of the same size as the separation sets. Recall that the leaves of the junction tree are associated with constraints in the optimal transport problem. Thus, the messages sent from these leaves correspond to an update step similar to the one in Algorithm 1. The messages sent within the graph encode the computation of the projections. More precisely, picking one leaf $c$ and sending the messages from all other leaves across the separation sets towards our leaf amounts to one exact computation of the projection $P_{c}({\bf K}\odot{\bf U}\odot{\bf M})$. However, the algorithm converges with any message schedule (Haasler et al. 2021c, Fan et al. 2022a).
Note that in the case where all cost are pair-wise and constraints act on one-marginals of the transport tensors, Algorithm 2 boils down to the Iterative Scaling Belief Propagation algorithms in Haasler et al. (2021c), Fan et al. (2022a). Here, we have extend on these methods and provide a more general algorithm for the case where we may have constraints on joint distributions, and where cost terms can depend on more than two marginals.
The Iterative Scaling Belief Propagation algorithm utilizes the graphical structure in problem 22-23 to significantly improve on the complexity of the Sinkhorn iterations. Recall that the computational bottleneck of Algorithm 1 lies in the computation of the projections, which require $\mathcal{O}(d^{J})$ operations. Algorithm 2 reduces this complexity by utilizing the graphical structure for computing the projections. The complexity of the problem depends then on the size of the messages, i.e., the size of the separation sets. This quantity can also be described by the so-called tree-width.
Definition 2 (Tree width (Bertele and Brioschi 1973, Halin 1976)).
In order for Algorithm 2 to be efficient we thus would like to pick a junction tree factorization with small tree-width. Given a graph, a junction tree decomposition that minimizes the tree-width is also called a minimal junction tree. In general, finding a junction tree with minimal width is a NP hard problem (Koller and Friedman 2009). However, for all applications in this paper it is easy to find minimal junction trees.
Example 2.
The junction trees in Figures 2, 3, and 4 are minimal junction trees for the respective graphs. Since the cliques in the former two junction trees are of cardinality 3, the respective graphs’ tree-widths are 2. Similarly, the last graph’s tree width is 3.
For Algorithm 2, we can directly apply the complexity analysis in Fan et al. (2022a). We will discuss the complexity result in remark 2.
@@ 4 Tree-structured multi-marginal optimal transport with local regularization
$\min_{\begin{subarray}{c}B_{(j,k)},(j,k)\in E\\ \mu_{j},j\in V\setminus\Gamma\end{subarray}}\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}\qquad\text{subject to}\quad B_{(j,k)}\mathbf{1}=\mu_{j},~{}~{}B_{(j,k)}^{\top}\mathbf{1}=\mu_{k}~{}~{}\text{for }(j,k)\in E,$ (28)
where $E$ is a set of all the edges of the graph. The matrices $C_{(j,k)}\in\mathbb{R}^{d\times d}$ and $B_{(j,k)}\in\mathbb{R}^{d\times d}$ represent the transport cost and transport plan corresponding to the edge connecting nodes $j$ and $k$, respectively. We implicitly assume the convention $B_{(j,k)}=B_{(k,j)}^{\top}$ and $C_{(j,k)}=C_{(k,j)}^{\top}$. In contrast to the global entropy regularization presented in Sec. 3 for a standard MOT problem 22, applying local regularization on each edge appears as a more direct and natural approach within this framework. Notably, in tree-structured optimal transport scenarios, such as barycenter problems, this form of regularization is frequently employed, as described in existing literature (Kroshnin et al. 2019).
@@ 4.1 Algorithm
Drawing inspiration from Kroshnin et al. (2019), we developed a parallel algorithm to address the coupled bi-marginal OT problem for general trees. Herein, we propose to use a modified entropy regularization similar to Fan et al. (2022a). The reference joint distribution between nodes $j$ and $k$ is denoted as:
$M_{(j,k)}:=\gamma_{j}\otimes\gamma_{k},$ (29)
@@
$\min_{\begin{subarray}{c}B_{(j,k)},(j,k)\in E\\ \mu_{j},j\in V\setminus\Gamma\end{subarray}}$ $\sum_{(j,k)\in E}\langle C_{(j,k)},B_{(j,k)}\rangle+\epsilon{\mathcal{H}}(B_{(j,k)}|M_{(j,k)})$ (30)
@@
The regularization, as formulated, is applied locally to each edge of the tree, in contrast to the previous formulations 22, which regularizes the multi-marginal optimal transport tensor. Utilizing a standard Lagrangian dual theory enables the derivation of the dual problem. Denote
$B_{(j,k)}=\operatorname{diag}(u_{(j,k)}\odot\gamma_{j})K_{(j,k)}\operatorname{diag}(u_{(k,j)}\odot\gamma_{k}),\qquad K_{(j,k)}=\exp(-C_{(j,k)}/\epsilon)$ (31)
where $u_{(j,k)}\in{\mathbb{R}}^{d}$ Furthermore, $u=\{u_{(j,k)}\}_{j\neq k}$ and $\rho=\{\rho_{j}\}_{j\in V\backslash\Gamma}$. For a node $j\in\Gamma$, its unique neighbor is denoted by $k_{j}$. The dual problem is then formulated as:
$\min_{u,\rho}~{}~{}f(u,\rho):=$ $\sum_{(j,k)\in E}\|B_{(j,k)}\|_{1}-\sum_{j\in\Gamma}\log(u_{(j,k_{j})})^{\top}\mu_{j}-\sum_{j\in V\backslash\Gamma}\rho_{j}$ (32)
@@
We then introduce another set of dual variables and apply the Lagrangian dual theory again to the dual problem 32, and obtain the update rules. The dual variables we introduce are reflected in $v_{j}$ in 34.
Theorem 1.
@@
$u_{(j,k)}=e^{\rho_{j}/|N(j)|}v_{j}./(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})),\qquad v_{j}=\bigg{(}\bigodot_{k\in N(j)}\Big{(}K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})\Big{)}\bigg{)}^{1/|N(j)|}.$ (34)
@@
Based on Theorem 1, we note that for any node $j$, the updates of $\{u_{(j,k)}\}_{k\in N(j)}$ or $\rho_{j}$ depend only on the variables $u_{(k,j)}$ from neighboring nodes $k\in N(j)$. This observation is critical for enabling parallel updates in the algorithm, as it allows for independently updating dual variables associated with two distinct parts of the tree. The concept of a bipartite graph becomes instrumental here, providing the framework for such parallelism.
Definition 3.
(Bipartite Graph (Metcalf and Casey 2016)). A bipartite graph is a graph where the vertices can be divided into two disjoint sets such that all edges connect a vertex in one set to a vertex in another set. There are no edges between vertices in the disjoint sets.
@@
[…]
To leverage this, the algorithm unfolds in the following manner: We begin by representing the tree as a bipartite graph, a natural representation since all trees are inherently bipartite according to Scheinerman (2012). The algorithm then iteratively updates until convergence, alternating between: i) Updating all dual variables from the first partition to the second. ii) Updating all dual variables from the second partition to the first. Denote the intermediate computational variables
$q^{t}_{(j,k)}=B^{t}_{(j,k)}\mathbf{1},\qquad q^{t+1}_{j}=\bigg{(}\bigodot_{k\in N(j)}q^{t}_{(j,k)}\bigg{)}^{\frac{1}{|N(j)|}}.$ (35)
Our method is detailed in Algorithm 3, termed Bipartite Iterative Scaling. A supporting proposition 1 underpins its theoretical foundation.
@@
Initialization of variables: $t=1,~{}~{}u^{0}_{(j,k)}=\mathbf{1}\in{\mathbb{R}}^{d}$ for $j\neq k$.
@@
       Let $S=S_{1}$ if $t$ is odd, otherwise $S=S_{2}$
      Calculate $q_{\{j,k\}}^{t},~{}~{}q_{j}^{t+1}$ according to 35 for $j\in S$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!u^{t+1}_{(j,k_{j})}=u^{t}_{(j,k_{j})}\odot\mu_{j}./q^{t}_{(j,k_{j})}\quad\text{for }j\in\Gamma\bigcap S$ (36)
$\!\!\!\!\!\!\!\!\!\!\!\!u_{(j,k)}^{t+1}=u_{(j,k)}^{t}\odot\frac{q_{j}^{t+1}}{\|q_{j}^{t+1}\|_{1}}./q_{(j,k)}^{t}\quad\text{for }j\in(V\backslash\Gamma)\bigcap S$ (37)
@@
            Round $B_{(j,k_{j})}$ by Algorithm 2 in Altschuler et al. (2017) such that $\widehat{B}_{(j,k_{j})}\in\Pi(\mu_{j},B^{\top}_{(j,k_{j})}{\bf 1})$

### v2

@@ 2.2.1 Generalized Euler flow
One of the first applications of multi-marginal optimal transport, the generalized Euler flow [BenCarCut15, Section 4.3], falls in fact into the category of a graph-structured problem. The Euler flow studies a relaxation of the incompressible Euler equation, where one seeks trajectories of particles between constrained initial and final data, that minimize the kinetic energy [brenier1989least]. Discretizing the problem into $J$ time points, we describe the flow by a $J$-mode multi-marginal optimal transport plan ${\mathbf{B}}\in{\mathbb{R}}_{+}^{d^{J}}$, where each marginal describes the density of the flow at time instance $j$, discretized into $d$ points. Note that the density of an incompressible flow, discretized uniformly over a given space, is described by a uniform distribution. Thus, we constrain each marginal of the transport plan by $P_{j}({\mathbf{B}})=\mu_{j}=\frac{1}{d}\bm{1}$. In the Euler flow, it is enforced that each particle at time 0 is transported to a given location at time $T$, given by a permutation $\sigma:\{1,\dots,d\}\to\{1,\dots,d\}$. This is modeled through a bi-marginal constraint $P_{1,J}({\mathbf{B}})=\bm{\mu}_{\{1,J\}}=\Pi_{\sigma}$, where $\Pi_{\sigma}\in{\mathbb{R}}^{d\times d}$ is the permutation matrix for the permutation $\sigma$. The Euler flow problem is thus formulated as the MOT problem
(10) $\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}$ $\langle{\bf C},{\mathbf{B}}\rangle$
subject to $P_{j}({\mathbf{B}})=\mu_{j}\text{ ~~for }j=1,\dots,J;\quad P_{1,J}({\mathbf{B}})=\bm{\mu}_{\{1,J\}},$
@@
[BenCarCut15] also propose an alternative relaxed formulation, where the bi-marginal constraint in Eq. 10 is replaced by an additional cost term in Eq. 11, and the problem becomes
(12) $\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j}\text{ for }j=1,\dots,J,$
@@
Both problems Eq. 10-Eq. 11 and Eq. 12-Eq. 13 can be described by a cyclic graph as illustrated in Fig. 1. In particular, we note that although the authors in [BenCarCut15] present only the computational method for the relaxed formulation Eq. 12-Eq. 13, our proposed framework encompasses both formulations.
Figure 1: Illustration of graphical structures in some applications. Gray nodes correspond to fixed marginals, and white nodes are estimated in the problem. The dotted edge between $x_{1}$ and $x_{J}$ in (a) may represent either a bi-marginal constraint, as in Eq. 10-Eq. 11, or a cost interaction, as in Eq. 12-Eq. 13.
(a) Euler flow
@@
(b) Wasserstein least squares
@@ 2.2.2 Wasserstein least squares
(15) $\min_{\pi,\mu_{0},\mu_{1}}$ $\int c(x_{0},x_{1},\cdots,x_{J},x_{J+1}){\rm{d}}\pi(x_{0},x_{1},\cdots,x_{J},x_{J+1})$
subject to $P_{j}(\pi)=\rho_{t_{j}},~~j=1,2,\cdots,J;\qquad P_{0}(\pi)=\mu_{0},\qquad P_{J+1}(\pi)=\mu_{1}.$
By discretizing the integral in Eq. 15 to a summation, we can transform the problem into the form presented in Eq. 9. To enforce that the optimality of the coupling between $\mu_{0}$ and $\mu_{1}$, we can add a large penalty on it, leading to the modified cost
(16) $c(x_{0},x_{1},\cdots,x_{J},x_{J+1})=\sum_{j=1}^{J}\|x_{j}-(1-t_{j})x_{0}-t_{i}x_{J+1}\|^{2}+\alpha\|x_{0}-x_{J+1}\|^{2}$
@@ 2.2.3 Wasserstein Splines
The Wasserstein spline developed in [CheConGeo18] is a natural notion of cubic spline over the space of probability distributions. Given a set of points $x_{t_{0}},x_{t_{2}},\ldots,x_{t_{J}}$ with $0=t_{0}<t_{1}<\cdots<t_{J}=1$, a cubic spline is a piece-wise cubic polynomial $\{x(t)\,\mid\,0\leq t\leq 1\}$ that is smooth and matches the given data points in the sense that $x(t_{j})=x_{t_{j}}$ for all $j=0,1,\ldots,J$. The cubic spline can also be characterized as the interpolating curve of the data points with minimum mean square acceleration. Based on this characterization, [CheConGeo18] generalized the notion of cubic spline to interpolate a set of probability distributions $\rho_{t_{0}},\rho_{t_{1}},\ldots,\rho_{t_{J}}$. The definition has a static reformulation in the phase space as
$\min_{\pi}\int c(x_{0},v_{0},\cdots,x_{J},v_{J}){\rm{d}}\pi(x_{0},v_{0},\cdots,x_{J},v_{J})\quad\text{ s.t. }P_{x_{j}}(\pi)=\rho_{t_{j}},\ j=0,1,\ldots,J,$
@@
$\sum_{i=0}^{J-1}\frac{1}{t_{\!j+1\!}-t_{j}}\!\left(\!12\left\|\frac{x_{\!j+1\!}-x_{j}}{t_{\!j+1\!}-t_{j}}-v_{j}\right\|^{2}\!\!-12\left\langle\frac{x_{\!j+1\!}-x_{j}}{t_{\!j+1\!}-t_{j}}-v_{j},v_{\!j+1\!}-v_{j}\right\rangle+4\|v_{\!j+1\!}-v_{j}\|^{2}\!\right)\!\!.$
@@
2.3 Multi-marginal optimal transport with global regulatization
One approach for solving multi-marginal optimal transport problems with a graph-structured cost Eq. 8 builds on regularizing the problem with a global entropic regularization term as in Eq. 4 [HaaSinZha20, fan2021complexity, HaaRinChe20, SinHaaZha20, ringh2023mean]. In this section we briefly present the core idea of this approach, since the parallel algorithms that we develop in the following sections builds on similar concepts. For a deeper introduction to the topic we refer the reader to the supplementary material, where we present the framework concisely, and in a more general setting than previous works.
It turns out that in case the underlying graph is a tree, i.e., it does not have any loops, computing a projection $P_{j}({\bf K}\odot{\bf U}\odot{\bf M})$, as needed in Algorithm 1, scales linearly in the number of marginals [HaaSinZha20], whereas a brute force approach for computing the projections scales exponentially. This result can be extended to general graphical models with possible loops, by partitioning the graph as a so-called junction tree [fan2021complexity]. A junction tree (also called tree decomposition) describes a partitioning of a graph, where several nodes are clustered together, such that the interactions between the clusters can be described by a tree structure.
Definition 2.2 (Junction tree [dawid1992applications]).
@@ 2.3 Multi-marginal optimal transport with global regulatization
A graph typically has several possible junction trees. In particular, one trivial junction tree is always given by a single cluster containing all nodes. For the computational method in [fan2021complexity] the junction tree is set up such that each constraint marginal or clique is represented by a leaf in the junction tree. The entropy-regularized graph structured optimal transport problem with junction tree decomposition $\mathcal{T}=(\mathcal{C},\mathcal{E})$ is thus of the form
(17) $\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}{|{\bf M}})\qquad\text{ subject to }P_{c}({\mathbf{B}})=\bm{\mu}_{c},\text{ for }c\in\mathcal{C}_{\Gamma},$
@@
(18) ${\bf C}({\bf x})=\sum_{c\in\mathcal{C}}{\bf C}_{c}({\bf x}_{c}).$
Such a junction tree representation exists for any graph, however it may require adding artificial clusters to the junction tree decomposition. Junction trees for the examples in Sections 2.2.12.2.3 are sketched in the supplementary material.
@@ 3 Tree-structured multi-marginal optimal transport with local regularization
(19) $\min_{\begin{subarray}{c}B_{(j,k)},(j,k)\in E\\ \mu_{j},j\in V\setminus\Gamma\end{subarray}}\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}\quad\text{ s.t. }B_{(j,k)}\mathbf{1}=\mu_{j},~~B_{(j,k)}^{\top}\mathbf{1}=\mu_{k}~~\text{for }(j,k)\in E,$
where $E$ is a set of all the edges of the graph. The matrices $C_{(j,k)}\in\mathbb{R}^{d\times d}$ and $B_{(j,k)}\in\mathbb{R}^{d\times d}$ represent the transport cost and transport plan corresponding to the edge connecting nodes $j$ and $k$, respectively. We implicitly assume the convention $B_{(j,k)}=B_{(k,j)}^{\top}$ and $C_{(j,k)}=C_{(k,j)}^{\top}$. In contrast to the global entropy regularization presented in Section 10 for a standard MOT problem Eq. 55, applying local regularization on each edge appears as a more direct and natural approach within this framework. Notably, in tree-structured optimal transport scenarios, such as barycenter problems, this form of regularization is frequently employed, as described in existing literature [kroshnin2019complexity].
@@ 3.1 Algorithm
Drawing inspiration from [kroshnin2019complexity], we developed a parallel algorithm to address the coupled bi-marginal OT problem for general trees. Herein, we propose to use a modified entropy regularization similar to [fan2021complexity]. The reference joint distribution between nodes $j$ and $k$ is denoted as:
(20) $M_{(j,k)}:=\gamma_{j}\otimes\gamma_{k},$
@@
(21) $\min_{\begin{subarray}{c}B_{(j,k)},(j,k)\in E\\ \mu_{j},j\in V\setminus\Gamma\end{subarray}}$ $\sum_{(j,k)\in E}\langle C_{(j,k)},B_{(j,k)}\rangle+\epsilon{\mathcal{H}}(B_{(j,k)}|M_{(j,k)})$
@@
The regularization, as formulated, is applied locally to each edge of the tree, in contrast to the previous formulations Eq. 55, which regularizes the multi-marginal optimal transport tensor. Utilizing a standard Lagrangian dual theory enables the derivation of the dual problem. Denote
(22) $B_{(j,k)}=\operatorname{diag}(u_{(j,k)}\odot\gamma_{j})K_{(j,k)}\operatorname{diag}(u_{(k,j)}\odot\gamma_{k}),\qquad K_{(j,k)}=\exp(-C_{(j,k)}/\epsilon)$
where $u_{(j,k)}\in{\mathbb{R}}^{d}$ . Furthermore, $u=\{u_{(j,k)}\}_{j\neq k}$ and $\rho=\{\rho_{j}\}_{j\in V\backslash\Gamma}$. For a node $j\in\Gamma$, its unique neighbor is denoted by $k_{j}$. The dual problem is then formulated as:
(23) $\min_{u,\rho}~~f(u,\rho)$= $\sum_{(j,k)\in E}\|B_{(j,k)}\|_{1}-\sum_{j\in\Gamma}\log(u_{(j,k_{j})})^{\top}\mu_{j}-\sum_{j\in V\backslash\Gamma}\rho_{j}$
@@
We then introduce another set of dual variables and apply the Lagrangian dual theory again to the dual problem Eq. 23, and obtain the update rules. The dual variables we introduce are reflected in $v_{j}$ in Eq. 25.
Theorem 3.1.
@@
(25) $u_{(j,k)}=e^{\rho_{j}/|N(j)|}v_{j}./(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})),\quad v_{j}\!=\bigg(\bigodot_{k\in N(j)}\!\Big(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})\Big)\bigg)^{\!1/|N(j)|}\!$
@@
Based on Theorem 3.1, we note that for any node $j$, the updates of $\{u_{(j,k)}\}_{k\in N(j)}$ or $\rho_{j}$ depend only on the variables $u_{(k,j)}$ from neighboring nodes $k\in N(j)$. This observation is critical for enabling parallel updates in the algorithm, as it allows for independently updating dual variables associated with two distinct parts of the tree. The concept of a bipartite graph becomes instrumental here, providing the framework for such parallelism.
Definition 3.2.
(Bipartite Graph [metcalf2016cybersecurity]). A bipartite graph is a graph where the vertices can be divided into two disjoint sets such that all edges connect a vertex in one set to a vertex in another set. There are no edges between vertices in the disjoint sets.
@@
$S_{1}$ $S_{2}$
To leverage this, the algorithm unfolds in the following manner: We begin by representing the tree as a bipartite graph, a natural representation since all trees are inherently bipartite according to [scheinerman2012mathematics]. The algorithm then iteratively updates until convergence, alternating between: i) Updating all dual variables from the first partition to the second. ii) Updating all dual variables from the second partition to the first. Denote the intermediate computational variables
(26) $q^{t}_{(j,k)}=B^{t}_{(j,k)}\mathbf{1},\qquad q^{t+1}_{j}=\bigg(\bigodot_{k\in N(j)}q^{t}_{(j,k)}\bigg)^{\frac{1}{|N(j)|}}.$
Our method is detailed in Algorithm 2, termed Bipartite Iterative Scaling. A supporting Proposition 3.3 underpins its theoretical foundation.
@@
Initialization of variables: $t=1,~~u^{0}_{(j,k)}=\mathbf{1}\in{\mathbb{R}}^{d}$ for $j\neq k$
@@
  0.2cm Let $S=S_{1}$ if $t$ is odd, otherwise $S=S_{2}$
  Calculate $q_{\{j,k\}}^{t},~~q_{j}^{t+1}$ according to Eq. 26 for $j\in S$
  -5pt
(27) $u^{t+1}_{(j,k_{j})}=u^{t}_{(j,k_{j})}\odot\mu_{j}./q^{t}_{(j,k_{j})}\qquad\text{for }j\in\Gamma\bigcap S$
  -10pt
(28) $u_{(j,k)}^{t+1}=u_{(j,k)}^{t}\odot\frac{q_{j}^{t+1}}{\|q_{j}^{t+1}\|_{1}}./q_{(j,k)}^{t}\qquad\text{for }j\in(V\backslash\Gamma)\bigcap S$
  -10pt
@@
     Round $B_{(j,k_{j})}$ by [altschuler2017near, Algorithm 2] such that $\widehat{B}_{(j,k_{j})}\in\Pi(\mu_{j},B^{\top}_{(j,k_{j})}{\bf 1})$ .